# Robust Point Matching with Distance Profiles

### v5

@@ 4.2 Statistical Complexity of the Third Lower Bound
where $C$ is an absolute constant that is independent of $n$, $d$, and $\mu$.
@@
where the last inequality follows from the Cauchy-Schwarz inequality. For any coupling $\gamma\in\Pi(\mu,\hat{\mu})$, note that $(\Phi\times\Phi)_{\#}\gamma\in\Pi(\Phi_{\#}\mu,\Phi_{\#}\hat{\mu})$, where $\Phi\times\Phi\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{t}\times\mathbb{R}^{t}$ is a map defined by $(\Phi\times\Phi)(x,x^{\prime})=(\Phi(x),\Phi(x^{\prime}))$. Conversely, for any coupling $\rho\in\Pi(\Phi_{\#}\mu,\Phi_{\#}\hat{\mu})$, we can find a coupling $\gamma\in\Pi(\mu,\hat{\mu})$ such that $\rho=(\Phi\times\Phi)_{\#}\gamma$ by Lemma 3.12 of [3]. In summary,
$\Pi(\Phi_{\#}\mu,\Phi_{\#}\hat{\mu})=\{(\Phi\times\Phi)_{\#}\gamma:\gamma\in\Pi(\mu,\hat{\mu})\}.$ (4.9)
Therefore,
$\begin{split}\inf_{\gamma\in\Pi(\mu,\hat{\mu})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}W_{1}(\mu_{x},\mu_{x^{\prime}})\,\mathrm{d}\gamma(x,x^{\prime})&\leq\inf_{\gamma\in\Pi(\mu,\hat{\mu})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\Phi(x)-\Phi(x^{\prime})\|_{2}\,\mathrm{d}\gamma(x,x^{\prime})\\ &=\inf_{\gamma\in\Pi(\mu,\hat{\mu})}\int_{\mathbb{R}^{t}\times\mathbb{R}^{t}}\|z-z^{\prime}\|_{2}\,\mathrm{d}(\Phi\times\Phi)_{\#}\gamma(z,z^{\prime})\\ &=\inf_{\rho\in\Pi(\Phi_{\#}\mu,\Phi_{\#}\hat{\mu})}\int_{\mathbb{R}^{t}\times\mathbb{R}^{t}}\|z-z^{\prime}\|_{2}\,\mathrm{d}\rho(z,z^{\prime})\\ &=W_{1}(\Phi_{\#}\mu,\Phi_{\#}\hat{\mu}),\end{split}$
where the first equality follows from the change of variables, and the second equality is due to (4.9). As $\Phi_{\#}\hat{\mu}$ is equivalent to the empirical measure of $\Phi_{\#}\mu$ which is defined on $\mathbb{R}^{t}$, we conclude from the standard results [14, 16] that $\mathbb{E}(*)\leq W_{1}(\Phi_{\#}\mu,\Phi_{\#}\hat{\mu})\leq O(n^{-1/t})$. ∎
We have used (4.5) to bound (**) defined in (4.4) by $\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})$. For each fixed $x\in\mathcal{X}$, as $\mu_{x}$ is a probability measure on $\mathbb{R}$, one can deduce from the standard results [14, 16] that $W_{1}(\mu_{x},\hat{\mu}_{x})$ enjoys a fast rate of convergence in terms of the sample size $n$, which is $O(n^{-1/2})$. Theorem 4 shows that the rate of convergence of $\mathbb{E}\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})$ is as fast as that of $\mathbb{E}W_{1}(\mu_{x},\hat{\mu}_{x})$ for fixed $x\in\mathcal{X}$, that is, $O(n^{-1/2})$. Theorem 4 considers the Euclidean case, which can be extended to general metric measure spaces; we provide this extension in Appendix A.3.
Theorem 4.
where $C$ is an absolute constant that is independent of $n$, $d$, and $\mu$.
Proof.
where $\mathcal{F}_{\mathcal{X}}:=\bigcup_{x\in\mathcal{X}}\mathcal{F}_{x}=\{\phi_{x,r}:x\in\mathcal{X},r\geq 0\}$. By Theorem 8.23 of [38],
@@
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})\right]\leq\mathbb{E}\left[\sup_{\phi\in\mathcal{F}_{\mathcal{X}}}|\mu\phi-\hat{\mu}\phi|\right]\leq C\sqrt{\frac{\mathrm{vc}(\mathcal{F}_{\mathcal{X}})}{n}}.$ (4.11)
where $C$ is an absolute constant that is independent of $x$, $n$, $\mathcal{X}$, and $\mu$. Here, $\mathrm{vc}(\mathcal{F}_{\mathcal{X}})$ is the Vapnik-Chervonenkis (VC) dimension of the function class $\mathcal{F}_{\mathcal{X}}$. As $\mathcal{X}\subset\mathbb{R}^{d}$, this is upper bounded by the VC dimension of the set of all closed Euclidean balls in $\mathbb{R}^{d}$, which is $d+1$; see [15]. Hence, we obtain (4.10). ∎
Lastly, we provide the analytic property of the distance between distance profiles used throughout this section.
Proposition 2.
Proof.
5 Simulations
5.1 Gaussian Mixture Matching
We first provide a simulation study to demonstrate the performance of the matching procedure discussed in Section 2. Figure 3(a) shows the probability of the perfect matching, namely, the left-hand side of (2.2). Here, $d=3$ and $K=10$, and we consider the clean case $t=s=K$ and the case with one outlier $t=s=K+1$ for the sample size $n=m\in\{500,1000\}$; also, we let $p_{k}=q_{k}=1/t$ and $\sigma=\sigma_{k}=\tau_{k}$ for all $k\in[t]$, and vary $\sigma$. We can observe that the probability of the perfect matching decreases as $\sigma$ grows for both cases. Though the presence of an outlier deteriorates the matching performance as expected, we can verify the outlier robustness of the proposed method in that the recovery probability is sufficiently high when $\sigma$ is small. Comparing the plots for $n=m=500$ and $n=m=1000$, we can see that the matching performance is better for the larger sample size as it better approximates the population-level distance profiles.
Although the analysis in Section 2 was based on the probability of the perfect matching, it is reasonable to expect matching errors in practice. Figure 3(b) shows the matching accuracy for the same setting as Figure 3(a) with $n=m=1000$, where the matching accuracy is the proportion of correctly matched $X_{i}$’s from the first $K$ components:
where $t,s$ are defined as in Theorem 1. Again, we can see that the proposed method performs reasonably well. Figure 3(c) compares the case with one outlier ($K=10$ and $t=s=11$) to the case with more outliers but with the same outlier proportion ($K=20$ and $t=s=22$) and to the case with more outliers with larger outlier proportions ($K=20$ and $t=s=24$). We can observe that the matching accuracy decreases as the outlier proportion increases, which is consistent with the theoretical results in Section 2. Even when the outlier proportion $(t-K)/K$ is the same, the matching accuracy is lower for larger $t$ as the larger effective dimensions of the underlying distributions $\mu,\nu$ lead to larger estimation errors given the same sample size. Figure 3(d) compares the case with $d=3$ to the case with $d=10$ for $K=10$ and $t=s=11$. We can see that the matching accuracy is lower and deteriorates much faster for larger $d$.
(a) Perfect matching probability
(b) Matching accuracy ($K=10$)
(c) Matching accuracy (varying $K$)
(d) Matching accuracy (varying $d$)
Figure 3: Gaussian mixture matching results. (a) shows the probability of the perfect matching shown on the left-hand side of (2.2) with $d=3$, $K=10$, and the sample size $n=m\in\{500,1000\}$; we let $p_{k}=q_{k}=1/t$ and $\sigma=\sigma_{k}=\tau_{k}$ for all $k\in[t]$, and vary $\sigma$. (b) shows the matching accuracy defined in (5.1) for the same setting as (a) with $n=m=1000$. (c) and (d) compare the matching accuracy for different settings with varying $K$ or $d$, while we keep the same sample size $n=m=1000$. For all plots, the sampling from $\mu,\nu$ is repeated 100 times. For (a), the marked points show the probability estimated from the 100 realizations. For (b)-(d), the marked points show the average of the matching accuracy over 100 realizations, while the shaded regions show the standard deviations. The blue solid lines of (b)-(d) are the same, which correspond to the case with $K=10$ and $t=s=11$.
5.2 Noise Stability
Next, we provide simulation results demonstrating the noise stability discussed in Section 3. We first obtain $n$ locations $\theta_{1},\ldots,\theta_{n}\in\mathbb{R}^{d}$ and define the other locations $\eta_{1},\ldots,\eta_{n}$ as in (3.2) using some rigid transformation $T$ and a permutation $\pi^{\ast}$. Then, we generate the points according to (3.3) and obtain $\hat{\pi}$ following the procedure explained in Theorem 2; we estimate $\mathbb{P}(\hat{\pi}=\pi^{\ast})$ by repeating this for 100 times. Figure 4 shows the results, where we vary $\sigma$—assuming $\sigma=\sigma_{i}=\tau_{i}$ for all $i\in[n]$—and compare with the probability $\mathbb{P}(\hat{\pi}^{\mathrm{OT}}=\pi^{\ast})$ for $\hat{\pi}^{\mathrm{OT}}$ obtained by the procedure (3.7). As shown in Figure 4(a), when the rigid transformation is based on a rotation matrix that only affects the first two coordinates, both methods can recover the true permutation to some extent for small enough $\sigma$; in this case, as the rotation matrix is relatively close to the identity matrix, the standard optimal transport method (3.7) shows good performance. For a rotation matrix that changes all the coordinates, however, the standard optimal transport method fails to recover the true permutation as shown in Figure 4(b), while the distance profile matching method performs almost as in Figure 4(a).
(a) Changing two coordinates
(b) Changing all coordinates
Figure 4: Noise stability results with $d=10$ and $n=100$. The plots show the probability $\mathbb{P}(\hat{\pi}=\pi^{\ast})$, where $\hat{\pi}$ corresponds to the one defined in Theorem 2 for the blue lines (distance profile) and to (3.7) for the red lines (standard OT); here, the probability is estimated by repeating each setting 100 times. For (a), the rigid transformation is based on a rotation matrix that only affects the first two coordinates, while (b) is based on a random rotation matrix changing all $d$ coordinates.
6 Conclusion
We have proposed a matching procedure based on distance profiles, which is easily implementable without having to solve nontrivial optimization problems over permutations. Most importantly, we have shown that the proposed method is provably robust to outliers and noise under some model assumptions that we believe capture essential aspects of empirical applications. To the best of our knowledge, this is the first work that provides theoretical guarantees for the robustness of a matching procedure based on distance profiles in the context of object matching. Also, we have extended the concept of distance profiles to the abstract setting and connected the proposed matching procedure to the Gromov-Wasserstein distance and its lower bound, with a new sample complexity result derived based on the properties of distance profiles. We have also demonstrated the empirical performance of the proposed method through real data and simulations.
We suggest several directions for future research. On the theoretical side, it would be interesting to consider other types of matching guarantees in Theorem 1. As mentioned in Section 5, the matching accuracy (5.1) may be practically more useful. Similarly, for Theorem 2, it would be interesting to consider the probability that $\hat{\pi}$ is close to $\pi^{\ast}$ under a certain discrepancy, such as the Hamming distance, instead of the probability of $\hat{\pi}=\pi^{\ast}$. On the practical side, it would be interesting to delve deeper into the procedure of finding inliers, that is, the points from the common part of two objects. In Remark 2, though we have noted that the threshold $\rho$ can be chosen to identify the inliers at least in theory, developing a practical method applicable to real data is very challenging in general. Lastly, building on the sample complexity result in Section 4, we may analyze the outlier robustness of the third lower bound in general cases. We leave these directions for future research.
Acknowledgments
Appendix A Appendix
A.1 Details of Remark 2
@@
First, recall from the proof of Theorem 1 that the event $C=(2R(Z_{1}+Z_{2})+2(U+V)<\Omega)$ implies $s(\pi(i))=t(i)$ for any $i\in[n]$ such that $t(i)\in[K]$. Hence, on the event $C$, one can deduce that for such $i\in[n]$, we have
@@
Now, let $C^{\prime}$ be the event $(2R(Z_{1}+Z_{2})+2(U+V)<\Omega^{\prime})$, where $\Omega^{\prime}$ represents the quantity in (2.10). Clearly, we have $\Omega^{\prime}\leq\Omega$ and thus $C^{\prime}\subset C$. Therefore, on the event $C^{\prime}$, we have
@@
Now, we have $\mathbb{P}(C^{\prime})\geq 1-\delta$ by repeating the proof of Theorem 1 with the left-hand side of (2.3), which is $\Omega$, replaced by $\Omega^{\prime}$. Therefore, we have (2.11) with probability at least $1-\delta$.
A.2 Lemmas
We have used the following lemmas on the concentration of the multinomial distribution and the chi-squared distribution in the proofs of Theorems 1 and 2. The former is an application of Proposition A.6.6 of [36], while the latter is a standard result that can be deduced from the Chernoff bound; for instance, see Chapter 2 of [38].
Lemma 3.
Lemma 4.
Let $\chi^{2}(d)$ be the chi-squared random variable with $d$ degrees of freedom. Then,
$\mathbb{P}\left(\chi^{2}(d)\geq d+t\right)\leq\exp\left(-\frac{t(t\wedge d)}{8d}\right)\quad\forall t\geq 0.$
Accordingly, for any $t\geq 2d$,
$\mathbb{P}\left(\chi^{2}(d)\geq t\right)\leq\exp\left(-\frac{t}{16}\right).$
A.3 Extension of Theorem 4 to General Metric Measure Spaces
From (4.11) in the proof of Theorem 4, one can deduce that Theorem 4 can be extended to any metric measure space $(\mathcal{X},d_{\mathcal{X}},\mu)$ by replacing $d+1$ in (4.10) with the VC dimension of $\mathcal{F}_{\mathcal{X}}$ defined in the proof, which is the VC dimension of the set of all balls in $\mathcal{X}$. Unfortunately, controlling the VC dimension of the set of all balls in a general metric measure space is highly nontrivial. A more straightforward quantity is the $\varepsilon$-covering number $\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})$, namely, the smallest possible number of $\varepsilon$-balls to cover $\mathcal{X}$. The following theorem derives an upper bound using the covering number for a general bounded metric measure space $(\mathcal{X},d_{\mathcal{X}},\mu)$.
Theorem 5.
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})\right]\leq\frac{C}{\sqrt{n}}\int_{0}^{\Delta/4}\sqrt{\frac{\Delta}{\varepsilon}+\log\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})}\,\mathrm{d}\varepsilon,$ (A.1)
Proof.
By the Kantorovich-Rubinstein theorem (Theorem 1.14 of [39]), we have
[…]
where =$\mathcal{F}=\{f\circ d_{\mathcal{X},x}:f\in\mathrm{Lip}_{1},x\in\mathcal{X}\}$ which is a function class from $\mathcal{X}$ to $\mathbb{R}$. Here, $d_{\mathcal{X},x}\colon\mathcal{X}\to\mathbb{R}$ is defined by $d_{\mathcal{X},x}(\cdot)=d_{\mathcal{X}}(x,\cdot)$. Using symmetrization and Dudley’s chaining argument (see Chapter 8 of [38]), one can derive the following:
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})\right]=\mathbb{E}\left[\sup_{\phi\in\mathcal{F}}\left(\mu\phi-\hat{\mu}\phi\right)\right]\leq\frac{C}{\sqrt{n}}\int_{0}^{\Delta/2}\sqrt{\log\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})}\,\mathrm{d}\varepsilon,$ (A.2)
where $\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})$ denotes the $\varepsilon$-covering number of $\mathcal{F}$ with respect to $\|\cdot\|_{\mathcal{X}}$, namely, the sup norm defined by $\|\phi\|_{\mathcal{X}}=\sup_{x\in\mathcal{X}}|\phi(x)|$ for any $\phi\colon\mathcal{X}\to\mathbb{R}$. Here, we are using the fact that $\sup_{\phi\in\mathcal{F}}\|\phi\|_{\mathcal{X}}\leq\Delta$ since $\sup_{z\in[0,\Delta]}|f(z)|\leq\Delta$ for any $f\in\mathrm{Lip}_{1}$ by Lipschitzness. Now, noticing that $\mathcal{F}$ is a function class consisting of composite functions, we upper bound $\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})$ by the product of the $\varepsilon$-covering number of $\mathrm{Lip}_{1}$ with respect to the sup norm $\|\cdot\|_{\infty}$ on $[0,\Delta]$, that is, $\|f\|_{\infty}=\sup_{z\in[0,\Delta]}|f(z)|$, and the $\varepsilon$-covering number of $\mathcal{X}$ with respect to $d_{\mathcal{X}}$. Concretely, we claim
$\mathcal{N}(2\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})\leq\mathcal{N}(\varepsilon,\mathrm{Lip}_{1},\|\cdot\|_{\infty})\times\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}}).$ (A.3)
Hence, we obtain (A.3). The standard upper bound on the covering number of $\mathrm{Lip}_{1}$ (see Exercise 8.2.7 of [38]) tells
$\log\mathcal{N}(\varepsilon,\mathrm{Lip}_{1},\|\cdot\|_{\infty})\leq\frac{c\Delta}{\varepsilon},$ (A.4)
where $c$ is some absolute constant. Combining (A.2), (A.3), and (A.4), we obtain (A.1). ∎
Remark 3.
If one assumes that $\mathcal{X}$ is a compact subset of $\mathbb{R}^{d}$ and $d_{\mathcal{X}}$ is the standard Euclidean distance in Theorem 5, the bounds coincide with those in Theorem 4. To see this, it suffices to notice that $\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})$ scales as $(1+2/\varepsilon)^{d}$; see Section 4.2 of [38]. The key difference, however, is that Theorem 5 requires compactness of the support of $\mu$, while Theorem 4 does not.

### v6

@@ C.1 Extension of Theorem 1 to General Metric Measure Spaces
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})\right]\leq\frac{C}{\sqrt{n}}\int_{0}^{\Delta/4}\sqrt{\frac{\Delta}{\varepsilon}+\log\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})}\,\mathrm{d}\varepsilon,$ (C.1)
@@
By the Kantorovich-Rubinstein theorem, e.g., Theorem 1.14 of [villani_2003], we have
where $\mathrm{Lip}_{1}=\left\{f\colon[0,\Delta]\to\mathbb{R}\,|\,f(0)=0~\text{and}~f~\text{is}~1\text{-Lipschitz}\right\}$. Here, we have used the fact that the supports of $\mu_{x}$ and $\hat{\mu}_{x}$ are contained in the interval $0,\Delta]$ for any $x\in\mathcal{X}$. Hence,
where $\mathcal{F}=\{f\circ d_{\mathcal{X},x}:f\in\mathrm{Lip}_{1},x\in\mathcal{X}\}$ which is a function class from $\mathcal{X}$ to $\mathbb{R}$. Here, $d_{\mathcal{X},x}\colon\mathcal{X}\to\mathbb{R}$ is defined by $d_{\mathcal{X},x}(\cdot)=d_{\mathcal{X}}(x,\cdot)$. Using symmetrization and Dudley’s chaining argument (see Chapter 8 of [vershynin_2018]), one can derive the following:
@@
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})\right]=\mathbb{E}\left[\sup_{\phi\in\mathcal{F}}\left(\mu\phi-\hat{\mu}\phi\right)\right]\leq\frac{C}{\sqrt{n}}\int_{0}^{\Delta/2}\sqrt{\log\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})}\,\mathrm{d}\varepsilon,$ (C.2)
where $\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})$ denotes the $\varepsilon$-covering number of $\mathcal{F}$ with respect to $\|\cdot\|_{\mathcal{X}}$, namely, the sup norm defined by $\|\phi\|_{\mathcal{X}}=\sup_{x\in\mathcal{X}}|\phi(x)|$ for any $\phi\colon\mathcal{X}\to\mathbb{R}$. Here, we are using the fact that $\sup_{\phi\in\mathcal{F}}\|\phi\|_{\mathcal{X}}\leq\Delta$ since $\sup_{z\in[0,\Delta]}|f(z)|\leq\Delta$ for any $f\in\mathrm{Lip}_{1}$ by Lipschitzness. Now, noticing that $\mathcal{F}$ is a function class consisting of composite functions, we upper bound $\mathcal{N}(\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})$ by the product of the $\varepsilon$-covering number of $\mathrm{Lip}_{1}$ with respect to the sup norm $\|\cdot\|_{\infty}$ on $0,\Delta]$, that is, $\|f\|_{\infty}=\sup_{z\in[0,\Delta]}|f(z)|$, and the $\varepsilon$-covering number of $\mathcal{X}$ with respect to $d_{\mathcal{X}}$. Concretely, we claim
$\mathcal{N}(2\varepsilon,\mathcal{F},\|\cdot\|_{\mathcal{X}})\leq\mathcal{N}(\varepsilon,\mathrm{Lip}_{1},\|\cdot\|_{\infty})\times\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}}).$ (C.3)
Hence, we obtain (C.3). The standard upper bound on the covering number of $\mathrm{Lip}_{1}$ (see Exercise 8.2.7 of [vershynin_2018]) tells
$\log\mathcal{N}(\varepsilon,\mathrm{Lip}_{1},\|\cdot\|_{\infty})\leq\frac{c\Delta}{\varepsilon},$ (C.4)
where $c$ is some absolute constant. Combining (C.2), (C.3), and (C.4), we obtain (C.1). ∎
Remark 6.
If one assumes that $\mathcal{X}$ is a compact subset of $\mathbb{R}^{d}$ and $d_{\mathcal{X}}$ is the standard Euclidean distance in Theorem 5, the bounds coincide with those in Theorem 1. To see this, it suffices to notice that $\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})$ scales as $(1+2/\varepsilon)^{d}$; see Section 4.2 of [vershynin_2018]. The key difference, however, is that Theorem 5 requires compactness of the support of $\mu$, while Theorem 1 does not.
C.2 Extension of Theorem 3: Identifying Points from the Common Part
Section 3.2 focuses on the matching of the points from the common part, namely, the first $K$ components. The matching guarantee ensures that all $X_{i}$’s such that $t(i)\in[K]$ are correctly matched, while ignoring the points from the outlier components. If we assume a stronger condition by replacing the left-hand side of (3.4) with the following quantity
$min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)\wedge\min_{\begin{subarray}{c}\alpha\in[t]\backslash[K]\\ \beta\in[s]\end{subarray}}\overline{W}(\alpha,\beta)-R\cdot\bigg(1-\sum_{k\in[K]}p_{k}\bigg),$ (C.5)
we can show that
$\max_{\begin{subarray}{c}i\in[n]\\ t(i)\in[K]\end{subarray}}W(i,\pi(i))<\min_{\begin{subarray}{c}i\in[n]\\ t(i)\notin[K]\end{subarray}}W(i,\pi(i))$ (C.6)
holds with probability at least $1-\delta$. This means that if the threshold $\rho$ in Algorithm 1 is chosen between the two quantities in (C.6), the output $I$ of the algorithm exactly coincides with the set $\{i\in[n]:t(i)\in[K]\}$, thereby identifying the points from the common part.
@@ C.2 Extension of Theorem 3: Identifying Points from the Common Part
Let us prove (C.6). To this end, first recall from the proof of Theorem 3 that the event $C=(2R(Z_{1}+Z_{2})+2(U+V)<\Omega)$ implies $s(\pi(i))=t(i)$ for any $i\in[n]$ such that $t(i)\in[K]$. Hence, on the event $C$, one can deduce that for such $i\in[n]$, we have
@@
Now, let $C^{\prime}$ be the event $(2R(Z_{1}+Z_{2})+2(U+V)<\Omega^{\prime})$, where $\Omega^{\prime}$ represents the quantity in (C.5). Clearly, we have $\Omega^{\prime}\leq\Omega$ and thus $C^{\prime}\subset C$. Therefore, on the event $C^{\prime}$, we have
@@
Now, we have $\mathbb{P}(C^{\prime})\geq 1-\delta$ by repeating the proof of Theorem 3 with the left-hand side of (3.4), which is $\Omega$, replaced by $\Omega^{\prime}$. Therefore, we have (C.6) with probability at least $1-\delta$.